OpenStudy (anonymous):

write an equation of the sine function with the amplitude and period. amplitude: 0.5, period: 6pi

OpenStudy (zzr0ck3r):

2pi/6pi = 1/3 so we have (1/2)sin((1/3)x)

OpenStudy (anonymous):

thanks! what is the period for y=-3sin0.5(theta)

OpenStudy (zzr0ck3r):

how did I get the 1/3? I took the 2p and divided it by the period

OpenStudy (anonymous):

no its a completely diff question

OpenStudy (zzr0ck3r):

so (2pi)/x = 1/2 is what you need to solve.

OpenStudy (zzr0ck3r):

its the same question in reverse:)

OpenStudy (anonymous):

i got 4pi for the second ?